# Characterization of Gaussian Tensor Ensembles

### v1

@@ 2.2 Proof of the Main Theorem
$f(H)$ $=\exp\bigg{(}-\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{a}{\Gamma_{i_{1},\ldots,i_{p}}}(H^{(0)}_{i_{1},\ldots,i_{p}})^{2}-\sum_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}\frac{a}{\Gamma_{i_{1},\ldots,i_{p}}}(H^{(1)}_{i_{1},\ldots,i_{p}})^{2}$
$\hskip 142.26378pt+\sum_{i_{1}\leq\ldots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}}H^{(0)}_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}+c\bigg{)}$
$=\exp\left(-a\sum_{i_{1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}\bar{H}_{i_{1},\ldots,i_{p}}+b\sum_{i_{1},\ldots,i_{p/2}}H_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}+c\right)$
@@
$f(\mathrm{H})$ $=\exp\bigg{(}-\sum_{\epsilon\in\mathcal{E}}\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{a}{\Gamma_{i_{1},\ldots,i_{p}}}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2}-\sum_{\epsilon\notin\mathcal{E}}\sum_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}\frac{a}{\Gamma_{i_{1},\ldots,i_{p}}}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2}$ (B)
$\hskip 142.26378pt+\sum_{i_{1}\leq\ldots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}}Q^{(\epsilon)}_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}+c\bigg{)}$
@@
$\|Q_{i_{1},\ldots,i_{p}}\|^{2}_{\mathrm{F}}$ $=\sum_{\iota_{1},\ldots,\iota_{p}\in\{1,2\}}|(Q_{i_{1},\ldots,i_{p}})_{\iota_{1},\ldots,\iota_{p}}|^{2}$
$=\sum_{\iota_{1},\ldots,\iota_{p}\in\{1,2\}}\mathrm{Re}((Q_{i_{1},\ldots,i_{p}})_{\iota_{1},\ldots,\iota_{p}})^{2}+\mathrm{Im}((Q_{i_{1},\ldots,i_{p}})_{\iota_{1},\ldots,\iota_{p}})^{2}$
$=\sum_{\epsilon:n_{1}(\epsilon)+n_{3}(\epsilon)\equiv 0[2]}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2}+\sum_{\epsilon:n_{1}(\epsilon)+n_{3}(\epsilon)\equiv 1[2]}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2}$
$=\sum_{\epsilon}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2}$
@@
$\|\mathrm{H}\|^{2}_{\mathrm{F}}=\sum_{i_{1},\ldots,i_{p}}\|Q_{i_{1},\ldots,i_{p}}\|^{2}_{\mathrm{F}}=\sum_{i_{1},\ldots,i_{p}}\sum_{\epsilon}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})^{2},$
and we deduce immediately from Equation (B) the desired result.
@@ 2.3 Letac extension
In this Section, we discuss briefly about a Letac-type extension of the Theorem. We say that a real symmetric tensor $H$ is isotropic if $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the unit Frobenius sphere of $\mathcal{S}^{(p)}(N)$. In the case $p=2$, such matrices have been studied by Kopp and Miller in [9] as Spherical Matrix Ensembles, but we would prefer the term isotropy instead of sphericity which could refer to rotational invariance in the physics literature. The original result for vectors by Letac is the following.
Proposition 2 (Letac [11]).
@@
A random vector $H\in\mathcal{S}^{(p)}(N)$ with $p\geq 2,N\geq 2$ has independent coordinates up to symmetries and is isotropic if and only if it belongs to the GOTE$(0,\gamma)$, $\gamma\in\mathbb{R}^{*}_{+}$.
Proof of the Corollary.
Let $p\geq 2,N\geq 2$, $K:=\sum_{i_{1}=1}^{N}\sum_{i_{2}=i_{1}}^{N}\ldots\sum_{i_{p}=i_{p-1}}^{N}1$ and $\Phi:\mathcal{S}^{(p)}(N)\to\mathbb{R}^{K}$ be defined by
$\Phi(H)=\Bigr{(}\sqrt{\Gamma_{i_{1},\ldots,i_{p}}}H_{i_{1},\ldots,i_{p}},1\leq i_{1}\leq\ldots\leq i_{p}\leq N\Bigr{)}.$
@@
Then $\Phi$ is linear, bijective, and an isometry in the sense that $|\Phi(H)|=\|H\|_{\mathrm{F}}$ for all $H\in\mathcal{S}^{(p)}(N)$. Next, $\Phi(H)$ is a random vector of $\mathbb{R}^{K}$, $K\geq 3$, with independent coordinates since $\Phi$ is a tensor function. Moreover, since $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the unit Frobenius sphere of $\mathcal{S}^{(p)}(N)$, and since $\Phi$ is a linear isometry, it follows that $\Phi(H)/|\Phi(H)|=\Phi(H/\|H\|_{\mathrm{F}})$ is uniformly distributed on the unit Euclidean sphere of $\mathbb{R}^{K}$. Therefore, by Proposition 2, $\Phi(H)$ is Gaussian, and therefore $H$ is Gaussian since $\Phi$ is linear and bijective. ∎
@@
The question is now : can we deduce Maxwell Theorem from this result as it is the case for vectors when $N\geq 3$ Note that for $p$ even and for all $\alpha\in\mathbb{R}$, $\alpha\mathcal{I}$ is orthogonal invariant, but $\alpha\mathcal{I}/\|\alpha\mathcal{I}\|_{\mathrm{F}}=\mathcal{I}/\sqrt{|\mathfrak{P}|}$ is not uniformly distributed on the unit Frobenius sphere. Moreover, if the law of $H$ is orthogonal invariant, then for all $\alpha\in\mathbb{R}$, the law of $H+\alpha I$ is also orthogonal invariant. We conjecture that if $H$ is orthogonal invariant with independent entries, then $H^{\prime}:=S-\mathbb{E}(S_{1,\ldots,1})\mathcal{I}$ is isotropic. This would give the Theorem by applying Corollary 1 to $H^{\prime}$. Remark that in the matrix case, if $H$ is orthogonal invariant with independent entries, $\mathbb{E}(H_{ij})=0$ for all $1\leq i<j\leq n$ and the law of $H_{kk}$ does not depend on $k$ so $H^{\prime}$ is centered, orthogonal invariant with independent entries. Is it sufficient for isotropy ? We leave this question open to the interested reader.
3 Basis of invariants
@@
Before proving that they form a basis of the invariant polynomials in the entries of a tensor, we define properly the trace invariants. This name was first given by Gurau in [6]. Let $\mathcal{G}_{n}^{(p)}$ the set of all connected $p$-regular multigraphs with $n$ vertices, that is graphs where multiedges and self-loops are allowed and each vertex has multiplicity $p$, with a given order for the $p$ edges around each vertex. We denote $\partial v$ the $p$-tuple of neighboring edges of $v$. The edges are indexed by $i_{1},\ldots,i_{np/2}$. Note than $np$ must be even, otherwise there is no trace invariant of rank $n$. Then, for $H$ a tensor of order $p$, dimension $N$ and $G=(V,E)\in\sqcup_{n\geq 0}\mathcal{G}_{n}^{(p)}$, we define the trace invariant associated to $G$ as the polynomial
$\mathfrak{m}_{G}(H)=\sum_{1\leq i_{1},\ldots,i_{E}\leq N}\prod_{v=1}^{N}H_{i_{(\partial v)_{1},\ldots,(\partial v)_{p}}}.$
@@
In the hermitian and self-dual hermitian cases, we ask moreover that the edges of the multigraphs connect two halfedges of different parity, that is $(\partial v)_{2t-1}$ and $(\partial w)_{2k}$ for two vertices $v$ and $w$ no necessarily distinct. Now we claim that the trace invariant are a basis of all the invariant polynomials in both three cases. Note that exactly the same result holds for polynomials in the entries of a family of tensors with possibly different orders where a basis is given by tensor maps (also called tensor networks) with vertices of multiplicity matching with the order of a tensor of the family.
@@ 3.1 Orthogonal case
The proof in the orthogonal case is already properly given by Kunisky, Moore and Wein in [10, Appendix A]. They recall references where the result appeared [19, 5] but they give a proof more suitable to our context. We recall here their path. Let $H$ be a real symmetric tensor of order $p$, dimension $N$ and $R$ be a polynomial of degree $n$ in the entries of $H$. Then there exists a vector of coefficients $C\in(\mathbb{R}^{N^{p}})^{n}\simeq(\mathbb{R}^{N})^{\otimes np}$ such that
$R(H)=\sum_{i_{1},\ldots,i_{np}}\underbrace{H_{i_{1},\ldots,i_{p}}\ldots H_{i_{(n-1)p+1},\ldots,i_{np}}}_{=(H^{\otimes n})_{i_{1},\ldots,i_{np}}}C_{i_{1},\ldots,i_{np}}=:\langle H^{\otimes n},C\rangle.$
@@
$R(H)=R(U^{T}\cdot H)=\langle U^{T}\cdot H^{\otimes n},C\rangle=\langle H^{\otimes n},U\cdot C\rangle=\langle H^{\otimes n},(U^{\otimes np})^{\sigma}\star_{np}C\rangle,$
[…]
$R(H)=\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma}\star_{np}C\rangle.$
It is well known [18, 3] that the expectation over the orthogonal group can be written in terms of the Weingarten functions, that is
$(\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma})_{i_{1},\ldots,i_{2np}}=\sum_{\nu,\tau\in\mathrm{Match}(np)}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t},i_{np+\tau(t)}},$
where $\mathrm{Match}(np)\in S_{np}$ is the set of involutions without fixed point (perfect matchings) and $\mathrm{Wg}_{\nu,\tau}\in\mathbb{R}$ is the Weingarten coefficient associated to $(\nu,\tau)$. Now we get
$(\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma}\star_{np}C)_{i_{1},\ldots,i_{np}}=\sum_{\tau\in\mathrm{Match}(np)}\alpha(\tau)\prod_{t=1}^{np}\delta_{i_{t},i_{\tau(t)}},$
and then denoting $S_{\tau}$ the tensor associated to the matching $\tau$ with entries $(S_{\tau})_{i_{1},\ldots,i_{p}}=\prod_{t=1}^{np}\delta_{i_{t},i_{\tau(t)}}$, we have
$R(H)=\sum_{\tau\in\mathrm{Match}(np)}\alpha(\tau)\langle H^{\otimes n},S_{\tau}\rangle=\sum_{\tau\in\mathrm{Match}(np)}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
where $G(\tau)$ is the multigraph canonically associated to $\tau$, that is a vertex $v\in\{0,\ldots,n-1\}$ has neighboring edges $\partial v=(vp+1,\ldots,(v+1)p)$ and the halfedge $(\partial v)_{k}$ is matched with $\tau((\partial v)_{k})$. This concludes the proof.

### v2

@@ 2.2 Proof of the main theorem
[…]
$+\sum_{i_{1}\leq\dots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}}H^{(0)}_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}+c\bigg)$
$={}$ $\exp\left(-a\sum_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}\bar{H}_{i_{1},\dots,i_{p}}+b\sum_{i_{1},\dots,i_{p/2}}H_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}+c\right)$
@@
[…]
$+\sum_{i_{1}\leq\dots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}}Q^{(\epsilon)}_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}+c\bigg)$ (2.3)
@@
$\|Q_{i_{1},\dots,i_{p}}\|^{2}_{\mathrm{F}}$ $=\sum_{\iota_{1},\dots,\iota_{p}\in\{1,2\}}|(Q_{i_{1},\dots,i_{p}})_{\iota_{1},\dots,\iota_{p}}|^{2}$
$=\sum_{\iota_{1},\dots,\iota_{p}\in\{1,2\}}\mathrm{Re}((Q_{i_{1},\dots,i_{p}})_{\iota_{1},\dots,\iota_{p}})^{2}+\mathrm{Im}((Q_{i_{1},\dots,i_{p}})_{\iota_{1},\dots,\iota_{p}})^{2}$
$=\sum_{\epsilon:n_{1}(\epsilon)+n_{3}(\epsilon)\equiv 0[2]}\bigl(Q^{(\epsilon)}_{i_{1},\dots,i_{p}}\bigr)^{2}+\sum_{\epsilon:n_{1}(\epsilon)+n_{3}(\epsilon)\equiv 1[2]}\bigl(Q^{(\epsilon)}_{i_{1},\dots,i_{p}}\bigr)^{2}$
$=\sum_{\epsilon}\bigl(Q^{(\epsilon)}_{i_{1},\dots,i_{p}}\bigr)^{2},$
@@
$\|\mathrm{H}\|^{2}_{\mathrm{F}}=\sum_{i_{1},\dots,i_{p}}\|Q_{i_{1},\dots,i_{p}}\|^{2}_{\mathrm{F}}=\sum_{i_{1},\dots,i_{p}}\sum_{\epsilon}\bigl(Q^{(\epsilon)}_{i_{1},\dots,i_{p}}\bigr)^{2},$
and we deduce immediately from equation (2.3) the desired result.
@@ 2.3 Letac extension
In this subsection, we discuss briefly about a Letac-type extension of the theorem. We say that a real symmetric tensor $H$ is isotropic if $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the unit Frobenius sphere of $\mathcal{S}^{(p)}(N)$. In the case $p=2$, such matrices have been studied by Kopp and Miller in [9] as spherical matrix ensembles, but we would prefer the term isotropy instead of sphericity which could refer to rotational invariance in the physics literature. The original result for vectors by Letac is the following.
Proposition 2.3 (Letac [11]).
@@
A random vector $H\in\mathcal{S}^{(p)}(N)$ with $p\geq 2,N\geq 2$ has independent coordinates up to symmetries and is isotropic if and only if it belongs to the $\mathrm{GOTE}(0,\gamma)$, $\gamma\in\mathbb{R}^{*}_{+}$.
Proof.
Let $p\geq 2,N\geq 2$, $K:=\sum_{i_{1}=1}^{N}\sum_{i_{2}=i_{1}}^{N}\dots\sum_{i_{p}=i_{p-1}}^{N}1$ and $\Phi\colon\mathcal{S}^{(p)}(N)\to\mathbb{R}^{K}$ be defined by
$\Phi(H)=\Bigr(\sqrt{\Gamma_{i_{1},\dots,i_{p}}}H_{i_{1},\dots,i_{p}},1\leq i_{1}\leq\dots\leq i_{p}\leq N\Bigr).$
@@
Then $\Phi$ is linear, bijective, and an isometry in the sense that $|\Phi(H)|=\|H\|_{\mathrm{F}}$ for all $H\in\mathcal{S}^{(p)}(N)$}. Next, $\Phi(H)$ is a random vector of $\mathbb{R}^{K}$, $K\geq 3$, with independent coordinates since $\Phi$ is a tensor function. Moreover, since $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the unit Frobenius sphere of $\mathcal{S}^{(p)}(N)$, and since $\Phi$ is a linear isometry, it follows that $\Phi(H)/|\Phi(H)|=\Phi(H/\|H\|_{\mathrm{F}})$ is uniformly distributed on the unit Euclidean sphere of $\mathbb{R}^{K}$. Therefore, by Proposition 2.3, $\Phi(H)$ is Gaussian, and therefore $H$ is Gaussian since $\Phi$ is linear and bijective. ∎
@@
The question is now: Can we deduce Maxwell theorem from this result as it is the case for vectors when $N\geq 3$? Note that for $p$ even and for all $\alpha\in\mathbb{R}$, $\alpha\mathcal{I}$ is orthogonal invariant, but $\alpha\mathcal{I}/\|\alpha\mathcal{I}\|_{\mathrm{F}}=\mathcal{I}/\sqrt{|\mathfrak{P}|}$} is not uniformly distributed on the unit Frobenius sphere. Moreover, if the law of $H$ is orthogonal invariant, then for all $\alpha\in\mathbb{R}$, the law of $H+\alpha I$ is also orthogonal invariant. We conjecture that if $H$ is orthogonal invariant with independent entries, then $H^{\prime}:=S-\mathbb{E}(S_{1,\dots,1})\mathcal{I}$ is isotropic. This would give the theorem by applying Corollary 2.4 to $H^{\prime}$. Remark that in the matrix case, if $H$ is orthogonal invariant with independent entries, $\mathbb{E}(H_{ij})=0$ for all $1\leq i<j\leq n$} and the law of $H_{kk}$ does not depend on $k$ so $H^{\prime}$ is centered, orthogonal invariant with independent entries. Is it sufficient for isotropy ? We leave this question open to the interested reader.
3 Complete family of invariants
@@
We define properly the trace invariants and then prove that any invariant polynomial can be written as a linear combination of the trace invariants. This name was first given by Gurau in [6]. Let $\mathcal{G}_{n}^{(p)}$ the set of all connected $p$-regular multigraphs with $n$ vertices, that is, graphs where multiedges and self-loops are allowed and each vertex has multiplicity $p$, with a given order for the $p$ edges around each vertex. We denote $\partial v$ the $p$-tuple of neighboring edges of $v$. The edges are indexed by $i_{1},\dots,i_{np/2}$. Note than $np$ must be even, otherwise there is no trace invariant of rank $n$. Then, for $H$ a tensor of order $p$, dimension $N$ and $G=(V,E)\in\bigsqcup_{n\geq 0}\mathcal{G}_{n}^{(p)}$, we define the trace invariant associated to $G$ as the polynomial
$\mathfrak{m}_{G}(H)=\sum_{1\leq i_{1},\dots,i_{E}\leq N}\prod_{v=1}^{N}H_{i_{(\partial v)_{1},\dots,(\partial v)_{p}}}.$
In the hermitian and self-dual hermitian cases, we ask moreover that the edges of the multigraphs connect two halfedges of different parity, that is, $(\partial v)_{2t-1}$ and $(\partial w)_{2k}$ for two vertices $v$ and $w$ no necessarily distinct.
Example 3.1.
Let $H$ be a (symmetric or hermitian) tensor of order $4$ and dimension $N$. We give an explicit example of trace invariant associated to a $4$-regular multigraph.
uuvvww112233445566
Figure 2: For this graph, we have $\partial u=(1,2,3,4),\partial v=(5,3,2,1)$ and $\partial w=(4,5,6,6)$. We decorate the vertices $u$, $v$ and $w$ by $H$ and each edge by an index ($i_{k}$ for edge $k$). The associated trace invariant is given by $\sum_{1\leq i_{1},\dots,i_{6}\leq N}H_{i_{1},i_{2},i_{3},i_{4}}H_{i_{5},i_{3},i_{2},i_{1}}H_{i_{4},i_{5},i_{6},i_{6}}$.
@@
We claim that the trace invariant are a complete family of all the invariant polynomials in both three cases. Note that exactly the same result holds for polynomials in the entries of a family of tensors, with possibly different orders. A complete family is given by trace invariants associated to graphs having vertices with multiplicity matching with the order of a tensor of the family (these polynomials are also called tensor maps or tensor networks).
@@ 3.1 Orthogonal case
The proof in the orthogonal case is already properly given by Kunisky, Moore and Wein in [10, Appendix A]. They recall references where the result appeared [5, 19] but they give a proof more suitable to our context. We recall here their path. Let $H$ be a real symmetric tensor of order $p$, dimension $N$ and $R$ be a polynomial of degree $n$ in the entries of $H$. Then there exists a vector of coefficients $C\in(\mathbb{R}^{N^{p}})^{n}\simeq\bigl(\mathbb{R}^{N}\bigr)^{\otimes np}$ such that
$R(H)=\sum_{i_{1},\dots,i_{np}}\underbrace{H_{i_{1},\dots,i_{p}}\dots H_{i_{(n-1)p+1},\dots,i_{np}}}_{=(H^{\otimes n})_{i_{1},\dots,i_{np}}}C_{i_{1},\dots,i_{np}}=:\big\langle H^{\otimes n},C\big\rangle.$
@@
$R(H)=R\bigl(U^{\mathsf{T}}\cdot H\bigr)=\big\langle U^{\mathsf{T}}\cdot H^{\otimes n},C\big\rangle=\big\langle H^{\otimes n},U\cdot C\big\rangle=\big\langle H^{\otimes n},\bigl(U^{\otimes np}\bigr)^{\sigma}\star_{np}C\big\rangle,$
with $\sigma=(2\ \ np+1)(4\ \ np+3)\cdots(np\ \ 2np-1)$, so $(U^{\otimes np,\sigma})_{i_{1},\dots,i_{2np}}=\prod_{t=1}^{np}U_{i_{t},i_{np+t}}$ (we contract along the odd legs because each $U$ has one contracted leg in $U\cdot C$). We can now symmetrize $C$ by taking the expectation over the Haar measure, that is,
$R(H)=\big\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma}\star_{np}C\big\rangle.$
It is well known [3, 18] that the expectation over the orthogonal group can be written in terms of the Weingarten functions, that is,
$\bigl(\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma}\bigr)_{i_{1},\dots,i_{2np}}=\sum_{\nu,\tau\in\mathcal{M}(np)}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t},i_{np+\tau(t)}},$
where $\mathcal{M}(np)\in S_{np}$ is the set of involutions without fixed point (perfect matchings) and $\mathrm{Wg}_{\nu,\tau}\in\mathbb{R}$} is the Weingarten coefficient associated to $(\nu,\tau)$. Now we get
$\bigl(\mathbb{E}_{U\in\mathrm{O}(N)}U^{\otimes np,\sigma}\star_{np}C\bigr)_{i_{1},\dots,i_{np}}=\sum_{\tau\in\mathcal{M}(np)}\alpha(\tau)\prod_{t=1}^{np}\delta_{i_{t},i_{\tau(t)}},$
and denoting $S_{\tau}$ the tensor associated to the matching $\tau$ with entries $(S_{\tau})_{i_{1},\dots,i_{p}}=\prod_{t=1}^{np}\delta_{i_{t},i_{\tau(t)}}$, we have
$R(H)=\sum_{\tau\in\mathcal{M}(np)}\alpha(\tau)\big\langle H^{\otimes n},S_{\tau}\big\rangle=\sum_{\tau\in\mathcal{M}(np)}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
where $G(\tau)$ is the multigraph canonically associated to $\tau$, that is, a vertex $v\in\{0,\dots,n-1\}$ has neighboring edges $\partial v=(vp+1,\dots,(v+1)p)$ and the halfedge $(\partial v)_{k}$ is matched with $\tau((\partial v)_{k})$. This concludes the proof.